OpenStudy (anonymous):

Mrs. Smith bought a 3 3/4 pound roast for $3.75 per pound. What was the total cost? Show your work please

OpenStudy (anonymous):

Oh gosh ok i'll try. So 3 3/4 = 3 + 3/4 = 3 + 0.75 = 3.75 3.75 x $3.75 = $14.06 i think. ;~;

OpenStudy (anonymous):

Ah! I see where I went wrong now! Thank you very much!

OpenStudy (anonymous):

aaah you're welcome